OpenStudy (anonymous):

Write the equation 2x - 3y = 6 in slope-intercept form. y = 3/2x + 3 y = 2/3x + 3 y = 2/3x - 2

OpenStudy (anonymous):

Solve for y

OpenStudy (anonymous):

how?

OpenStudy (anonymous):

2x - 3y = 6 -3y = 6 - 2x 3y = 2x - 6 y = 2x / 3 - 2

OpenStudy (anonymous):

thanks for the explain/////xp

OpenStudy (anonymous):

2x - 3y = 6 \[2x -6= 3y \rightarrow 3y= 2x-6 \rightarrow y= \frac{2}{3}x -\frac{6}{3}\] i.e. \[y= \frac{2}{3}x -2\] which is in the frequired slope intercept form where slope =m= 2/3 and y-intercept=c=-2